OpenStudy (timinthor):

Please help! Solve for m m^-5=1/32

OpenStudy (igreen):

'm' is to the -5th power, do you know what to do to get rid of that exponent of -5?

OpenStudy (igreen):

\(\sf x^{-a} = \dfrac{1}{x^a}\)

OpenStudy (igreen):

Can you change \(\sf m^{-5}\) into that form?

OpenStudy (timinthor):

No, i dont know, normally what i do is just solve it in my head with the "guess and test" method but i think that m=2 but im not sure

OpenStudy (igreen):

Yes..2 is correct.

OpenStudy (igreen):

Here's another way to do it. \(\sf m^{-5} \rightarrow \dfrac{1}{m^5}= \dfrac{1}{32}\) Cross multiply: \(\sf \dfrac{1}{m^5}\nearrow \dfrac{1}{32}\) \(\sf \dfrac{1}{m^5}\searrow \dfrac{1}{32}\) \(\sf m^5 = 32\) To undo the exponent of 5, remember that \(\sf \sqrt{x^2} = x\). So find the 5th root of both sides. \(\sf \sqrt[5]{m^5} = \sqrt[5]{32}\) \(\sf m = \sqrt[5]{32}\) \(\sf m = 2\)

OpenStudy (igreen):

Get it?

OpenStudy (igreen):

The opposite of squaring is finding the square root, so the opposite of an exponent of 5 is finding the 5th root.

OpenStudy (igreen):

Guess-and-check is also an easy way to solve this problem.

OpenStudy (timinthor):

Ohh, that makes more sense, i dont really know why this problem is on my test, we havent been doing things with squaring but oh well :) thank you